Solve by elimination
5x + 5y = -7
7x - 2y = 19
If all work could be shown (even the simplest) it would be very appreciated.
I am having trouble figuring out how to simplify 80/10 down to 16/5.
Thank you!

Found 2 solutions by jorel1380, stanbon:
Answer by jorel1380(3719)

10x+10y=-14
35x-10y=95
-------------
45x=81
x=81/45
"=9/5
y=-16/5..

Answer by stanbon(75887)

Solve by elimination
5x + 5y = -7
7x - 2y = 19
-----------------------
Multiply thru the 1st by 2:
Multiply thru the 2nd be 5:
10x + 10y = -14
35x - 10y = 5*19
-----
Add and solve for "x":
45x = 81
---
x = 81/45
--
x = 9/5
-----
Solve for "y":
5x + 5y = -7
5(9/5)+5y = -7
9 + 5y = -7
5y = -16
y = -16/5
=================
Cheers,
Stan H.
==========